If Sam's age after 10 years will be triple times of his age before 6 years, how old is he right now?
Answer by dkppathak(439)

If Sam's age after 10 years will be triple times of his age before 6 years, how old is he right now?
let present age is X yrs
before 6 yrs age was =X-6
after 10 yrs now =X+10
as per given conditions
X+10=3(x-6)
x+10=3x-18
2x=28
x=14